OpenStudy (volleyballlover55):

The volume of a cone is 141.3 cubic inches. The height of the cone is 15 inches. What is the radius of the cone, rounded to the nearest inch? Use π = 3.14. 2 inches 3 inches 5 inches 9 inches

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (mayankdevnani):

write the formula of volume of cone first

OpenStudy (volleyballlover55):

ok

OpenStudy (volleyballlover55):

v=pi*r^2 h/3

OpenStudy (mayankdevnani):

now plug in the values

OpenStudy (volleyballlover55):

ok

OpenStudy (volleyballlover55):

v = 3.14 * 15^2 *3?

OpenStudy (mayankdevnani):

nope

OpenStudy (volleyballlover55):

?

OpenStudy (mayankdevnani):

we have to find radius

OpenStudy (volleyballlover55):

ok..

OpenStudy (mayankdevnani):

try again

OpenStudy (volleyballlover55):

can you help please with that part. i dont know it, i tried

OpenStudy (mayankdevnani):

\[\large 141.3=\frac{3.14 \times r^2 \times 15}{3}\]

OpenStudy (mayankdevnani):

solve for r

OpenStudy (volleyballlover55):

ok

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (volleyballlover55):

its 3,-3..

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (mayankdevnani):

good and radius can't be negative so, 3 is correct

OpenStudy (volleyballlover55):

ok awesome! next question

OpenStudy (mayankdevnani):

post it

OpenStudy (volleyballlover55):

@mayankdevnani

OpenStudy (mayankdevnani):

Ugh,use distance formula

OpenStudy (volleyballlover55):

which is...

OpenStudy (mayankdevnani):

lmao you forget

OpenStudy (volleyballlover55):

lol yes i do sorry

OpenStudy (mayankdevnani):

lol search it again

OpenStudy (volleyballlover55):

what do i search?

OpenStudy (mayankdevnani):

distance formula girl

OpenStudy (volleyballlover55):

ok

OpenStudy (volleyballlover55):

(x1,y1) and (x2,y2)

OpenStudy (volleyballlover55):

the problem is i dont know what to put in those..

OpenStudy (mayankdevnani):

what are the co ordinates of C and D?

OpenStudy (volleyballlover55):

2 and 6?

OpenStudy (mayankdevnani):

nope

OpenStudy (mayankdevnani):

i mean in terms of x-co ordinate and y-co ordinate

OpenStudy (volleyballlover55):

2 and -4?

OpenStudy (mayankdevnani):

C(0,2) and D(6,-4)

OpenStudy (mayankdevnani):

do you know how did i get it?

OpenStudy (volleyballlover55):

oohhhh yes i get it!

OpenStudy (mayankdevnani):

now apply distance formula

OpenStudy (volleyballlover55):

but i dont know what to do...

OpenStudy (volleyballlover55):

can you please write the formula and ill solve it..

OpenStudy (mayankdevnani):

\[\large \bf distance=\sqrt{(6-0)^2+(-4-2)^2}\]

OpenStudy (volleyballlover55):

thank you!

OpenStudy (mayankdevnani):

np

OpenStudy (volleyballlover55):

i got 8.4?

OpenStudy (mayankdevnani):

option B would be correct

OpenStudy (volleyballlover55):

ok next! What is the length, in units, of the hypotenuse of a right triangle if each of the two legs is 3 units? Square root of 18 units Square root of 27 units 3 units 6 units

OpenStudy (mayankdevnani):

draw the diagram first

OpenStudy (volleyballlover55):

what diagram?

OpenStudy (mayankdevnani):

diagram of triangle

OpenStudy (volleyballlover55):

i dont know how to do that

OpenStudy (mayankdevnani):